OpenStudy (anonymous):

Hello, in this question I am asked to "Evaluate the derivative at the given point"... "y=3x^2-4x; (-1,7)"

OpenStudy (bahrom7893):

y' = 6x - 4, at x = -1 y' = -6 - 4 = -10

OpenStudy (bahrom7893):

Please click on become a fan if I helped, I really want to get to the next level!! Thanks =)

OpenStudy (anonymous):

What did you do to come up with 6x-4?

OpenStudy (bahrom7893):

differentiate with respect to x